NEED HELP WITH THESE QUESTIONS PLEASE..VERY IMPORTANT THANKS
1. For the relation y=-3(x+5)²-4
a) Find the equation of the axis of symmetry
b) Find the direction of opening
-
2. Write the equation below in vertex form by completing the square.
y=2x²+12x-3
-
3. For the quadratic relation 2xV+x-15=0
a)determine the number of real roots
b) solve the equation
-
4. Describe the transformation applied to the graph of y=x² to obtain the graph of y=½(x+6)²+12

Answer by stanbon!(97)

1a. axis of symmetry x=-5
1b the -3 tells you the opening is down
2. y-3 =2(x^2+6x)
y-3+18 =2(x^2+6x+9)
y+15 =2(x+3)^2
y = 2(x+3)^2-15
3. Your posting is garbled.
Cheers,
Stan H.